You invest $3,500 for three years at 8 percent.
a. What is the value of your investment after one year? Multiply $3,500 × 1.08.

  
b. What is the value of your investment after two years? Multiply your answer to part a by 1.08. (Round your answer to 2 decimal places.)


  
c. What is the value of your investment after three years? Multiply your answer to part b by 1.08. This gives your final answer. (Round your answer to 2 decimal places.)
Answer #1

a.  Value of investment after one year = $3,500 * 1.08 = $3,780

b.  Value of investment after two year = $3,780 * 1.08 = $4,082.40

c. Value of investment after three year = $4,082.40 * 1.08 = $4,409.00